Can you construct a linear transformation T : R (3) 2 → R

3

such that

Im(T) = {(x, y,z) ∈ R

3

: ax + by + cz = 0} where a, b, c ∈ R are constants?


1
Expert's answer
2021-06-23T06:14:05-0400

The aim is to construct a transformation "T:\mathbb{R}^3\rightarrow\mathbb{R}^3" such that "Im(T)=\{(x,y,z)\in{\mathbb{R}}^3|ax+by+cz=0\}". We can set: "T(x,y,z)=(u,v,w)", where the point "(u,v,w)" is the orthogonal projection of "(x,y,z)" on the plane given by equation "ax+by+cz=0."